Investment X offers to pay you $3,902 per year for 9 years, whereas Investment Y offers to pay you $5,199 per year for 6 years. What is the dollar difference (higher minus lower) in the present values of these two cash flow streams if the discount rate is 7.32 percent? Answer to two decimals.
Answer #1

Present value of Investment X:-

=PV(rate,nper,pmt)

=PV(7.32%,9,3902)

=25079.95

Present value of Investment Y:-

=PV(7.32%,6,5199)

=24538.31

Difference in present value =25079.95-24538.31 =541.64

Where,

rate is rate of interest

nper is number of periods

pmt is annuity